# Better Language Models Exhibit Higher Visual Alignment

### v2

@@ 4 Language-driven Visual Generalization
5 LLMs for General-Purpose VLMs
The previous section revealed the intriguing capabilities of decoder-based LLMs to encode visually relevant information in their representations. Next, we investigate whether this inherent visual alignment can be leveraged to obtain effective and efficient multimodal models by integrating off-the-shelve LLMs directly into CLIP-like VLMs. For this, we relax the strict zero-shot setup and utilize larger-scale image-caption datasets to explore the benefits and limitations of general-purpose VLMs that fuse existing frozen foundation models with minimal data and compute.
5.1 Experimental Setup
Methodology.
We propose integrating LLMs into a CLIP-like framework through “Shared Vision-Language-Locked Tuning“ (ShareLock), extending the architecture and training setup introduced in Section 3. As illustrated in Figure 2, this design places the LLM at the core of the model and necessitates expressive language representations for strong VLM performance. Given the increased availability of training data compared to the strict zero-shot setting, we increase the projection network to four layers with a hidden dimension of . The additional parameters, combined with ReLU activations between layers, enhance ShareLock’s ability to capture complex, non-linear cross-modal correspondences. Additionally, the maximum number of steps is increased to 5000.
Datasets.
Our investigation focuses on leveraging LLMs with minimal additional paired data and explores how unimodal embeddings can drive robust multimodal performance with minimal supervision and alignment. As a result, our evaluation is limited to comparably small paired datasets.
COCO Captions. COCO Captions (Chen et al., 2015) contains 83k images with multiple human-written captions per image, from which we randomly sample during training. CC3M. Conceptual Captions (Sharma et al., 2018) comprises 2.8M filtered web-scraped image-alt-text pairs. We also utilize a smaller subset with more balanced concept coverage designed for LLaVA (Liu et al., 2023). CC12M. CC12M (Changpinyo et al., 2021) expands CC3M by 12M image-text pairs, enabling evaluation at larger scales. Due to expired links, our version contains about 8.5M samples.
Training speed and storage.
Precomputing features for the CC3M-Llava subset (563k pairs) using Llama-3 8B requires around 8 hours on a single A100 40GB GPU. Extracting DINOv2 features and optimizing the MLP-based projection network take 1 GPU each, totaling roughly 10 GPU hours. This is paired with a significant storage reduction from over 80GB for the raw data to just 12GB for precomputed features.
Evaluation.
We employ a comprehensive suite of evaluations to assess ShareLock’s capabilities across a wide range of tasks. Based on the publicly available CLIP Benchmark (LAION AI, 2022), we gauge the models’ zero-shot classification abilities across diverse datasets: ImageNet-1k (Deng et al., 2009), ImageNet-R(endition) (Hendrycks et al., 2021a), ImageNet-A(dversarial) (Hendrycks et al., 2021c), ImageNet-S(ketch) (Wang et al., 2019), Oxford-IIIT Pets (Parkhi et al., 2012), Oxford Flowers (Nilsback and Zisserman, 2008), Stanford Cars (Krause et al., 2013), and FGVCAircraft (Maji et al., 2013). We also provide qualitative text-to-image retrieval results on ImageNet for CC3M-trained models. Moreover, the challenging compositionality Winoground task (Thrush et al., 2022) is explored.
Baselines and comparisons.
We compare our straightforward and economical way of incorporating LLMs into VLMs to more conventional CLIP-like methods, particularly emphasizing data-efficient alignment approaches. Alongside the original ViT-B/16 variant of CLIP (Radford et al., 2021), we test against several CLIP-like models trained on public datasets of different scales and modified learning objectives (Fan et al., 2023; Mu et al., 2022; Gadre et al., 2023; Mu et al., 2022). Leveraging pretrained models, we evaluate how ShareLock compares to LiT (Zhai et al., 2022) and ASIF (Norelli et al., 2023) by reproducing these methods on smaller-scale datasets. For LiT baselines, we initialize the language encoder with pretrained BERT weights (Devlin et al., 2019), following Zhai et al. (2022). When comparing LiT, ASIF, and ShareLock models, the exact same precomputed input features (barring the language component of LiT) are used.
5.2 Comparison to Conventional VLMs
@@
The fine-grained nature of certain classification problems (cols. “Pet“ to “Aircraft“ in Tab. 2) demands larger-scale datasets with more diverse and nuanced concepts included as minute visual differences may be insufficiently captured in the text space. Consequently, low performance on small-scale datasets can be observed across all methods. Nonetheless, the LLM representations still contain visually valuable signals, enabling ShareLock to surpass alternative methods trained on the same data in 15/16 cases and demonstrating effective utilization of intrinsic LLM knowledge to generalize to truly novel concepts.
Multi-Lingual Understanding.
Most popular multimodal datasets consist primarily of English captions. Consequently, VLMs like CLIP and LiT experience significant performance drops when performing inference in other languages, as shown in Table 3 on multi-lingual ImageNet LAION AI (2022). In contrast, the broader and typically multi-lingual pretraining of LLMs enables ShareLock to harness cross-linguistic image-text consistencies, greatly mitigating the performance loss in non-English languages. Even with substantially fewer training samples, ShareLock surpasses the original CLIP model, achieving an accuracy of versus for Chinese and versus for Italian. In comparison, both LiT and CLIP models similarly trained on CC12M demonstrate near-random performance in these languages. This transfer of capabilities makes the inclusion of LLMs especially attractive for low-resource languages with little available or high-quality multimodal data.
@@
Figure 4 illustrates that ShareLock achieves comparable performance to CLIP and DataComp models using orders of magnitude less data. Utilizing frozen LLM features is especially effective in low-data regimes, consistently outperforming conventional CLIP-like models, further underlining their visually-relevant semantic content and capacity to facilitate generalization.
@@ 5.3 Qualitative Results
In addition to quantitative evaluations, Figure 3 demonstrates ShareLock’s strong text-image alignment across diverse prompts, showing particular advantages over CC3M-trained CLIP and LiT models for both fine-grained (e.g., “a photo of a BMW") and abstract (e.g., “[…] heavy seas") queries.
5.4 Choice of Language Model
As the nature and quality of the frozen language features are of great significance in the proposed architecture, we ablate the choice of language model on the CC3M dataset. Reflecting the insights from our investigation in Section 4, Table 5 highlights the potential of decoder-based models for vision-language tasks. Although BERT encoders serve as the starting point in LiT models, they perform poorly without fine-tuning. Similarly, all-roberta-v1 (Liu et al., 2019) improves significantly but remains inferior to LLM-based representations, despite being highlighted by Maniparambil et al. (2024) for its high inherent visual alignment. In contrast, frozen decoder-based representations consistently surpass BERT-based ones, with gains of to , showcasing the richness of strong LLM representations from text-only pretraining. Even naive last-token-pooling is suitable for extracting visual information from LLMs and the more sophisticated multi-token embedding approach of NV-Embed (Lee et al., 2024) fails to materialize in consistent and notable performance gains over off-the-shelf Llama and Gemma models.
Further analysis into the role of vision backbones and using more sophisticated projection network architectures can be found in Appendix D.
| Type | Language Model | IN1k | IN-R | IN-A | Pets | Cars | Aircraft |
| Enc. | BERT-Base | ||||||
| all-roberta-v1 | |||||||
| Dec. | Llama-3 8B | ||||||
| Gemma-2 9B | |||||||
@@ 6 Conclusion
7 Limitations
While we compare a wide range of publicly available encoder- and decoder-based language models, attributing concrete performance differences on mainly architectural differences is not possible due to the different pretraining objectives, training datasets, and number of parameters used in these models. As the primary focus of this work is on highlighting the visual alignment inherent to unimodal language models and their incorporation into data- and compute-efficient vision-language models, our investigation is limited to datasets with up to 12M image-caption pairs. Scaling ShareLock to larger datasets for further performance improvements is left for future work. Although pre-computing features only once is possible due to locked vision and language backbones, using LLMs with billions of parameters significantly increases computational costs of forward passes compared to conventional encoder architectures.
@@ 8 Acknowledgments
We gratefully acknowledge the ELLIS Unit Amsterdam for providing funding for a research visit to Copenhagen. We thank SURF for providing GPU cluster access and support during this project. We acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) with hardware funded by the German Research Foundation (DFG). This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1.
@@ Appendix A Reproducibility Statement
Appendix B Acquisition of Textual Class Representations
More details about the characteristics and acquisition of these class representations are provided in Section B of the appendix. Besides the template-based targets proposed by Radford et al. (2021) that solely substitute the respective class names, we generate more comprehensive auxiliary information about classes (e.g., Wikipedia-style articles) using the instruction-tuned version of the Llama-3-8B model and acquire human-curated information from Wikipedia (details provided in B).
@@ Appendix B Acquisition of Textual Class Representations
A set of 80 human-engineered prompt templates in the style of "a photo of a <class name>" are adopted from Radford et al. (2021).
@@
Appendix C In-Depth Analysis of Visual LLM Generalization
@@
As seen in Figure 5 and Table 12, the text encoder taken from the CLIP ViT-B/16 model constitutes a strong baseline that is not yet reached by any of the current state-of-the-art LLMs. However, considering the explicit vision alignment on 400M image-text pairs makes the strong generalization abilities unsurprising. On the contrary, obtaining a score of as the best performing unimodal model, Gemma-2 9B (Mesnard et al., 2024) is close to CLIP’s despite no multimodal exposure. This further underscores the remarkable degree of visual alignment inherent to decoder-based language models.
LLMs are often subject to additional task-specific fine-tuning. Table 6 compares different models and their derivatives tuned on instruction and multimodal data. While vision-tuned variants are available for Phi-3 (Microsoft, 2024), Qwen2 (Wang et al., 2024a), and Vicuna v1.5 (Zheng et al., 2023), XTuner’s LLaVA model (XTuner Contributors, 2023) constitutes the source for the visual Llama-3 variant. Across all models, with the exception of Llama-3, the impact of fine-tuning is minor, typically shifting performance by only a few decimal points in the case of Phi-3 and Vicuna. Interestingly, Llama-3’s performance declines significantly after fine-tuning ( and for instruction and visual tuning), contrasting with the generally stable results of other models. Neither instruction-based nor visual fine-tuning shows a clear and consistent advantage in improving overall performance. These insights are also reflected in Figure 5 and Table 12. Ultimately, the base model’s architecture and training regime are more significant in determining performance than post-hoc fine-tuning strategies.
@@
Appendix D Additional Ablations
D.1 Projection Network Architecture
The multi-layer perceptron (MLP) projection networks of ShareLock as introduced in Section 3 are conceivably simple. As these are the only unfrozen and tunable parts of the model architecture and thus responsible for aligning vision and language inputs, they are of particular significance to aptly process and transform the inputs. Following Zhai et al. (2022), no transformation to the vision inputs is applied for any of the architectures. With a hidden size of 4096 and four layers, the MLP processing the language features comprises approximately 53M parameters.
In addition to the straightforward MLP-based networks, also more sophisticated Transformer-based architectures are inspired by recent works. First introduced as part of the BLIP-2 model (Li et al., 2023), the Q-Former is a lightweight Transformer-based model that extracts features from an input modality using cross-attention with learnable query tokens. Similarly, albeit introduced in a different context, NV-Embed (Lee et al., 2024) uses a latent attention layer to pool language tokens and receive a global embedding. Slight adjustments are made to both baseline architectures to better suit late-fusion vision-language modeling, which we will denote as Q-Transformer and NV-Transformer. While features are extracted via last-token-pooling (see Figure 6) when used with MLP projection networks, all tokens of the input sequence are considered for the alternative architectures. The hyperparameters were selected based on the implementation details suggested in the original publications and to approximately match the MLP baseline in learnable parameter count. Both the Q-Transformer and the NV-Transformer projection networks have a token dimension of 1024 in the Transformer parts of the models, eight learnable queries (Q-Transformer), and key/values (NV-Transformer). Whereas the Q-Transformer consists of 3 blocks and 4 attention heads, NV-Transformer comprises a total of four layers with eight cross-attention heads each.
The choice of projection network architecture is ablated in Table 7. The evaluated models use DINOv2-ViT-B/14 as their vision backbone. While no single architecture consistently scores best, the MLP-based ShareLock configuration performs competitively compared to NV-Transformer and Q-Transformer throughout the evaluation datasets. Additionally, Transformer-based architectures entail increased computational complexity due to the more evolved attention mechanism and processing of more tokens, making MLPs an attractive choice from an efficiency perspective as well. These results suggest that the additional information contained across all tokens of an input is not significantly more adjuvant compared to solely considering the last token representation as is done with the MLP.
@@ D.1 Projection Network Architecture
D.2 Loss Function
The use of the Sigmoid Loss (SigLIP) has unlocked further efficiency and performance enhancements in the standard CLIP regime (Zhai et al., 2023). However, no substantial gains are found when using SigLIP in the ShareLock framework as presented in Table 8.
@@ D.2 Loss Function
D.3 Choice of Vision Encoder
| Language Model | IN1k | IN-R | IN-A | Pets | Cars | Aircraft |
| ResNet101 (sup.) | ||||||
| ConvNextV2 (sup.) | ||||||
| ViT-L (sup.) | ||||||
| DINOv2-L (self-sup.) |
As ShareLock is agnostic to the utilized vision encoder, we ablate variants with differing architectures and supervision regimes in Table 9. Since language embeddings map to the vision space, encoder choice is a crucial factor, as reflected in notable performance differences. Unlike DINOv2, the other models were only trained on ImageNet and exhibit lower robustness and generality, highlighting the benefits of broad pretraining across diverse concepts – even without explicit supervision.
One advantage of CLIP is its favorable scaling characteristics when increasing the size of the vision encoder (Radford et al., 2021). To validate if comparable trends are present in ShareLock, we vary DINOv2-based backbones ranging from Small to Giant vision transformers and present the results in Table 10. Indeed, ShareLock also profits from scaling up the vision backbones with the average scores increasing by and , when moving from the Small to the Base and from the Base to the Large DINOv2 models, respectively. However, the benefits of scale start to level off thereafter, and only marginal differences are present when utilizing representations from the Giant vision encoder.
| Variant | IN1k | IN-R | IN-A | Pets | Cars | ESAT |
| Small | ||||||
| Base | ||||||
| Large | ||||||
| Giant |
Appendix E Extended Visual Generalization Results
Due to limited space, the reported results on visual generalization in Table 1 and Figure 1 of the main paper are averaged accuracies across five seeds and four datasets (AWA2 (Xian et al., 2017), CUB (Wah et al., 2011), FGVCAircraft (Maji et al., 2013), and ImageNet+). For increased transparency, the results are presented without dataset-level aggregation in Tables 11 and 12.
| Class Representation | Type | Language Model | AWA2 | CUB | Aircraft | ImageNet+ | Average |
| Class Names | Decoder | Llama-3 8B | |||||

### v3

@@ 5.2 Comparison to Conventional VLMs
The fine-grained nature of certain classification problems (cols. “Pet“ to “Aircraft“ in Tab. 2) demands larger-scale datasets with more diverse and nuanced concepts included as minute visual differences may be insufficiently captured in the text space. Consequently, low performance on small datasets can be observed across all methods. Nonetheless, the LLM representations still contain visually valuable signals, enabling ShareLock to surpass other methods trained on the same data in 15/16 cases and demonstrating effective utilization of intrinsic LLM knowledge to generalize to truly novel concepts.
Multilingual Understanding.
Most popular multimodal datasets consist primarily of English captions. Consequently, VLMs like CLIP and LiT experience significant performance drops when performing inference in other languages, as shown in Table 3 on multilingual ImageNet LAION AI (2022). In contrast, the broader and typically multilingual pretraining of LLMs enables ShareLock to harness cross-linguistic image-text consistencies, greatly mitigating the performance loss in non-English languages. Even with substantially fewer training samples, ShareLock surpasses the original CLIP model, achieving an accuracy of versus for Chinese and versus for Japanese. In comparison, both LiT and CLIP models similarly trained on CC12M demonstrate near-random performance in these languages. This transfer of capabilities makes the inclusion of LLMs especially attractive for low-resource languages with little available or high-quality multimodal data.
@@
| Method | Language Backbone | |
| BERT-base | Llama-3 (8B) | |
| CLIP | ||
| LiT | ||
| ShareLock | ||
Figure 5 illustrates that ShareLock achieves comparable performance to CLIP and DataComp models while using orders of magnitude less data. Utilizing frozen LLM features is especially effective in low-data regimes, consistently outperforming conventional CLIP-like models, further underlining their visually-relevant semantic content and capacity to facilitate generalization. Additionally, tuning far fewer parameters enables substantially larger batch sizes (see Table 5), which has been shown to improve contrastive learning performance (Zhai et al., 2022). Frozen backbones also enables the integration of large-scale LLMs into CLIP-like architectures while maintaining training efficiency.
@@ 5.3 Qualitative Results
In addition to quantitative evaluations, Figure 6 demonstrates ShareLock’s strong text-image alignment across diverse prompts, showing advantages over CC3M-trained CLIP and LiT models for both fine-grained (e.g., “a photo of a BMW") and abstract (e.g., “[…] heavy seas") queries.
5.4 Component Analysis
Choice of language model
As the nature and quality of the frozen language features are of great significance in the proposed architecture, we examine the choice of language model on the CC3M dataset. Reflecting the insights from our investigation in Section 4, Table 6(a) highlights the potential of decoder-based models for vision-language tasks. Although BERT encoders serve as the starting point in LiT models, they perform poorly without fine-tuning. Similarly, all-roberta-v1 (Liu et al., 2019) improves significantly but remains inferior to LLM-based representations, despite being highlighted by Maniparambil et al. (2024) for its high inherent visual alignment. In contrast, frozen decoder-based representations consistently surpass BERT-based ones, with gains of to , showcasing the richness of strong LLM representations from the more extensive pretraining and larger parameter counts. Even naive last-token-pooling is suitable for extracting visual information from LLMs and the more sophisticated multi-token embedding approach of NV-Embed (Lee et al., 2024) fails to materialize in consistent and notable performance gains over off-the-shelf Llama and Gemma models.
| Lang. Model | IN1k | IN-R | IN-A | Pets | Cars | |
| Enc. | BERT-Base | |||||
| all-roberta-v1 | ||||||
| Dec. | Llama-3 8B | |||||
| Gemma-2 9B | ||||||
@@ 5.4 Component Analysis
| Vision Model | IN1k | IN-R | IN-A | Pets | Cars | |
| Sup. | ResNet101 | |||||
| ConvNextV2-L | ||||||
| ViT-L | ||||||
| Self-sup. | DINOv2-S | |||||
| -B | ||||||
| -L | ||||||
| -G |
Choice of vision encoder
As ShareLock is agnostic to the utilized vision encoder, we compare variants with differing architectures and supervision regimes in Table 6(b). Since language embeddings map to the vision space, encoder choice is a crucial factor, as reflected in notable performance differences. Unlike DINOv2, the other models were only trained on ImageNet and exhibit lower robustness and generality on other datasets, highlighting the benefits of broad pretraining across diverse concepts – even without explicit supervision.
One advantage of CLIP is its favorable scaling characteristics when increasing the size of the vision encoder (Radford et al., 2021). To validate if comparable trends are present in ShareLock, we vary DINOv2-based backbones ranging from Small to Giant vision transformers (cf. Tab. 6(b)). Indeed, ShareLock also profits from scaling up the vision backbones with the average scores increasing by and , when moving from the Small to the Base and from the Base to the Large DINOv2 models, respectively. However, the benefits of scale start to level off thereafter, and only marginal differences are present when utilizing representations from the Giant vision encoder.
Choice of projection network depth.
As the only learnable parameters, the choice of the projection networks and is crucial for aligning the backbones and generalizing across downstream tasks. Table 7 reveals that projecting vision features into the language embedding space consistently underperforms compared to the reverse direction. This suggests that the vision embedding space is more semantically coherent and continuous, allowing the model to better approximate the latent concept manifold, particularly in low-data regimes. Conversely, language embeddings appear to encode non-semantic, visually irrelevant information (e.g., syntactic features), which are difficult to predict solely from visual inputs. Similarly, introducing projection layers atop both backbones and thereby removing the fixed reference space leads to reduced performance and signs of overfitting.
| IN1k | IN-R | IN-A | Pets | Cars | Aircraft | ||
| - | 2 | ||||||
| - | 4 | ||||||
| 2 | - | ||||||
| 4 | - | ||||||
| 6 | - | ||||||
| 2 | 2 | ||||||
| 4 | 4 |
@@ 7 Acknowledgments
The authors thankfully acknowledge the HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander Universität Erlangen-Nürnberg (FAU) under the BayernKI project v115be. BayernKI funding is provided by Bavarian state authorities. We thank SURF for providing GPU cluster access and support during this project. This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1 and TNO’s research programs Appl.AI and DSS. We gratefully acknowledge the ELLIS Unit Amsterdam for providing funding for a research visit to Copenhagen.
@@ Appendix A Reproducibility Statement
Appendix B Limitations
While we compare a wide range of publicly available encoder- and decoder-based language models, attributing concrete performance differences on mainly architectural differences is not possible due to the different pretraining objectives, training datasets, and number of parameters used in these models. As the primary focus of this work is on highlighting the visual alignment inherent to unimodal language models and their incorporation into data- and compute-efficient vision-language models, our investigation is limited to datasets with up to 12M image-caption pairs. Scaling ShareLock to larger datasets for further performance improvements is left for future work. Although pre-computing features only once is possible due to locked vision and language backbones, using LLMs with billions of parameters significantly increases computational costs of forward passes compared to smaller conventional encoder architectures.
Appendix C Acquisition of Textual Class Representations
Besides the template-based targets proposed by Radford et al. [2021] that solely substitute the respective class names, we generate more comprehensive auxiliary information about classes (e.g., Wikipedia-style articles) using the instruction-tuned version of the Llama-3-8B model and acquire human-curated information from Wikipedia (details provided in C).
@@ Appendix C Acquisition of Textual Class Representations
A set of 80 human-engineered prompt templates in the style of "a photo of a <class name>" are adopted from Radford et al. [2021].
@@
Appendix D In-Depth Analysis of Visual LLM Generalization
@@
As seen in Figure 7 and Table 12, the text encoder taken from the CLIP ViT-B/16 model constitutes a strong baseline that is not yet reached by any of the current state-of-the-art LLMs. However, considering the explicit vision alignment on 400M image-text pairs makes the strong generalization abilities unsurprising. On the contrary, obtaining a score of as the best performing unimodal model, Gemma-2 9B [Mesnard et al., 2024] is close to CLIP’s despite no multimodal exposure. This further underscores the remarkable degree of visual alignment inherent to decoder-based language models.
LLMs are often subject to additional task-specific fine-tuning. Table 8 compares different models and their derivatives tuned on instruction and multimodal data. While vision-tuned variants are available for Phi-3 [Microsoft, 2024], Qwen2 [Wang et al., 2024b], and Vicuna v1.5 [Zheng et al., 2023], XTuner’s LLaVA model [XTuner Contributors, 2023] constitutes the source for the visual Llama-3 variant. Across all models, with the exception of Llama-3, the impact of fine-tuning is minor, typically shifting performance by only a few decimal points in the case of Phi-3 and Vicuna. Interestingly, Llama-3’s performance declines significantly after fine-tuning ( and for instruction and visual tuning), contrasting with the generally stable results of other models. Neither instruction-based nor visual fine-tuning shows a clear and consistent advantage in improving overall performance. These insights are also reflected in Figure 7 and Table 12. Ultimately, the base model’s architecture and training regime are more significant in determining performance than post-hoc fine-tuning strategies.
@@
Appendix E Additional Ablations and Comparisons
E.1 Projection Network Architecture
The multi-layer perceptron (MLP) projection networks of ShareLock as introduced in Section 3 are conceivably simple. As these are the only unfrozen and tunable parts of the model architecture and thus responsible for aligning vision and language inputs, they are of particular significance to aptly process and transform the inputs. Following Zhai et al. [2022], no transformation to the vision inputs is applied for any of the architectures. With a hidden size of 4096 and four layers, the MLP processing the language features comprises approximately 53M parameters.
In addition to the straightforward MLP-based networks, also more sophisticated Transformer-based architectures are inspired by recent works. First introduced as part of the BLIP-2 model [Li et al., 2023], the Q-Former is a lightweight Transformer-based model that extracts features from an input modality using cross-attention with learnable query tokens. Similarly, albeit introduced in a different context, NV-Embed [Lee et al., 2024] uses a latent attention layer to pool language tokens and receive a global embedding. Slight adjustments are made to both baseline architectures to better suit late-fusion vision-language modeling, which we will denote as Q-Transformer and NV-Transformer. While features are extracted via last-token-pooling (see Figure 8) when used with MLP projection networks, all tokens of the input sequence are considered for the alternative architectures. The hyperparameters were selected based on the implementation details suggested in the original publications and to approximately match the MLP baseline in learnable parameter count. Both the Q-Transformer and the NV-Transformer projection networks have a token dimension of 1024 in the Transformer parts of the models, eight learnable queries (Q-Transformer), and key/values (NV-Transformer). Whereas the Q-Transformer consists of 3 blocks and 4 attention heads, NV-Transformer comprises a total of four layers with eight cross-attention heads each.
The choice of projection network architecture is compared in Table 9. The evaluated models use DINOv2-ViT-B/14 as their vision backbone. While no single architecture consistently scores best, the MLP-based ShareLock configuration performs competitively compared to NV-Transformer and Q-Transformer throughout the evaluation datasets. Additionally, Transformer-based architectures entail increased computational complexity due to the more evolved attention mechanism and processing of more tokens, making MLPs an attractive choice from an efficiency perspective as well. These results suggest that the additional information contained across all tokens of an input is not significantly more adjuvant compared to solely considering the last token representation as is done with the MLP.
@@ E.1 Projection Network Architecture
E.2 Loss Function
The use of the Sigmoid Loss (SigLIP) has unlocked further efficiency and performance enhancements in the standard CLIP regime [Zhai et al., 2023]. However, no substantial gains are found when using SigLIP in the ShareLock framework as presented in Table 10.
@@ E.2 Loss Function
Appendix F Extended Visual Generalization Results
Due to limited space, the reported results on visual generalization in Table 1 and Figure 1 of the main paper are averaged accuracies across five seeds and four datasets (AWA2 [Xian et al., 2017], CUB [Wah et al., 2011], FGVCAircraft [Maji et al., 2013], and ImageNet+). For increased transparency, the results are presented without dataset-level aggregation in Tables 11 and 12. Moreover, we include more detailed results corresponding to Figure 3 in Table 13. In addition to last token aggregation for decoder-based models, we also evaluate applying the same mean pooling across all tokens as utilized for encoders. However, last token pooling results in superior performance.
| Class Repr. | Type | Language Model | AWA2 | CUB | Aircraft | IN+ | Avg |
| Class Names | Dec | Llama-3 8B | |||||